# Constant Metric Scaling in Riemannian Computation

### v1

@@ Abstract
Constant rescaling of a Riemannian metric appears in many computational settings, often through a global scale parameter that is introduced either explicitly or implicitly. Although this operation is elementary, its consequences are not always made clear in practice and may be confused with changes in curvature, manifold structure, or coordinate representation. In this note we provide a short, self-contained account of constant metric scaling on arbitrary Riemannian manifolds. We distinguish between quantities that change under such a scaling, including norms, distances, volume elements, and gradient magnitudes, and geometric objects that remain invariant, such as the Levi–Civita connection, geodesics, exponential and logarithmic maps, and parallel transport. We also discuss implications for Riemannian optimization, where constant metric scaling can often be interpreted as a global rescaling of step sizes rather than a modification of the underlying geometry. The goal of this note is purely expository and is intended to clarify how a global metric scale parameter can be introduced in Riemannian computation without altering the geometric structures on which these methods rely.
@@ 1 Introduction
Riemannian geometry has become an increasingly common tool in modern computational settings, including optimization on manifolds, geometric data analysis, and representation learning (bhattacharya_2012_NonparametricInferenceManifolds; patrangenaru_2016_NonparametricStatisticsManifolds; boumal_2023_IntroductionOptimizationSmooth). In many of these applications, one encounters a global parameter that controls the overall scale of the Riemannian metric. Such a parameter is often introduced implicitly—for example through a temperature, normalization constant, or so-called curvature parameter—and is sometimes interpreted as modifying the geometric structure of the underlying space. While this intuition is natural, it can obscure a simple but important distinction: when a Riemannian metric is rescaled by a constant factor, the resulting transformation alters measurements but leaves the underlying geometry unchanged.
@@
The purpose of this note is to provide a clear and self-contained account of constant metric scaling on Riemannian manifolds and to spell out its consequences for Riemannian computation. The geometric facts themselves are classical and well known in differential geometry. However, their implications are not always made explicit in computational or modeling contexts, where metric scaling is frequently conflated with changes in curvature, manifold constraints, or coordinate representations. In particular, it is not uncommon for constant metric rescaling to be treated as if it modifies geodesics, exponential maps, or even the manifold itself. One motivation for this paper is to disentangle these ideas and to record, in one place, what does and does not change under constant metric scaling.
@@
We focus on a simple but practically relevant setting: given a Riemannian manifold $(M,g)$, we consider a rescaled metric $\tilde{g}=\lambda g$ with $\lambda>0$ constant. Under this transformation, numerical quantities such as norms, distances, volumes, and gradient magnitudes are rescaled in a predictable way. At the same time, the Levi–Civita connection, geodesics, exponential and logarithmic maps, and parallel transport remain exactly the same. As a consequence, many computational tools used in practice, particularly those based on geodesic constructions, are invariant under constant metric scaling and require no modification. From a practical view, the effect of metric scaling can often be interpreted as a global rescaling of step sizes rather than a change in geometric structure.
@@
This paper is intentionally expository. We do not aim to introduce new theory or algorithms, but rather to give explicit derivations and a systematic classification of invariant and variant quantities under constant metric scaling. The presentation is kept general, applying to arbitrary Riemannian manifolds, and is meant to serve as a compact reference for readers who wish to incorporate a flexible metric scale parameter into Riemannian models or algorithms without inadvertently altering the geometry they rely on. Throughout, we adopt a deliberately modest perspective: the results collected here may be viewed as a small exercise in clarifying familiar geometric facts, motivated by their recurring appearance in contemporary computational practice.

### v2

@@ Abstract
Constant rescaling of a Riemannian metric appears in many computational settings, often through a global scale parameter that is introduced either explicitly or implicitly. Although this operation is elementary, its consequences are not always made clear in practice and may be confused with changes in curvature, manifold structure, or coordinate representation. In this note, we provide a short, self-contained account of constant metric scaling on arbitrary Riemannian manifolds. We distinguish between quantities that change under such a scaling, including norms, distances, volume elements, and gradient magnitudes, and geometric objects that remain invariant, such as the Levi–Civita connection, geodesics, exponential and logarithmic maps, and parallel transport. We also discuss implications for Riemannian optimization, where constant metric scaling can be interpreted as a global rescaling of step sizes rather than a modification of the underlying geometry. The goal of this note is purely expository and is intended to clarify how a global metric scale parameter can be introduced in Riemannian computation without altering the geometric structures on which these methods rely.
@@ 1 Introduction
Riemannian geometry has become an increasingly common tool in modern computational settings, including optimization on manifolds (boumal_2023_IntroductionOptimizationSmooth) and geometric data analysis (bhattacharya_2012_NonparametricInferenceManifolds; patrangenaru_2016_NonparametricStatisticsManifolds). In many of these applications, one can encounter a global parameter that controls the overall scale of the Riemannian metric. Such a parameter is often introduced implicitly through a temperature, normalization constant, or so-called curvature parameter, and is sometimes interpreted as modifying the geometric structure of the underlying space. While this intuition is natural, it can obscure a simple but important distinction: when a Riemannian metric is rescaled by a constant factor, the resulting transformation alters measurements but leaves the underlying geometry unchanged.
@@
The purpose of this note is to provide a self-contained account of constant metric scaling on Riemannian manifolds and to spell out its consequences for Riemannian computation. The geometric facts themselves are classical and well known in differential geometry (docarmo_1992_RiemannianGeometry; lee_1997_RiemannianManifoldsIntroduction). However, their implications are not always made explicit in computational or modeling contexts, where metric scaling is frequently conflated with changes in curvature, manifold constraints, or coordinate representations. In particular, it is not uncommon for constant metric rescaling to be treated as if it modifies geodesics, exponential maps, or even the manifold itself. One motivation for this paper is to disentangle these ideas and to record what does and does not change under constant metric scaling in one place.
@@
We focus on a simple but practically relevant setting. Given a Riemannian manifold $(M,g)$, we consider a rescaled metric $\tilde{g}=\lambda g$ with $\lambda>0$ constant. Under this transformation, numerical quantities such as norms, distances, volumes, and gradient magnitudes are rescaled in a predictable way. At the same time, the Levi–Civita connection, geodesics, exponential and logarithmic maps, and parallel transport remain exactly the same. As a consequence, many computational tools used in practice, particularly those based on geodesic constructions, are invariant under constant metric scaling and require no modification. From a practical view, the effect of metric scaling can often be interpreted as a global rescaling of step sizes rather than a change in geometric structure.
@@
This paper is intentionally expository. We do not aim to introduce new theory or algorithms, but rather to give explicit derivations and a systematic classification of invariant and variant quantities under constant metric scaling. The presentation is kept general, applying to arbitrary Riemannian manifolds, and is meant to serve as a compact reference for readers who wish to incorporate a flexible metric scale parameter into Riemannian models or algorithms without inadvertently altering the geometry they rely on. Throughout, we adopt a deliberately modest perspective and the results collected here may be viewed as a small exercise in clarifying familiar geometric facts, motivated by their recurring appearance in contemporary computational practice.